OpenStudy (anonymous):

What is the distance from the point (6,0) to the line y=3x+2?

OpenStudy (amistre64):

which distance you want? shortest most likly

OpenStudy (amistre64):

the shortest distance is a perpendicular line....

OpenStudy (amistre64):

so to find it, we need the line that is perp to this and the point they intersect

OpenStudy (anonymous):

Possible answers are 4, \[2\sqrt{10} , \sqrt{26} , and\ 3\sqrt{6}\] so I guess which ever equals one of those

OpenStudy (amistre64):

the perp has a product of slopes that equals -1 3x = -1 x = -1/3 0 = -1/3(6) +b b = 2 y = (-1/3)x +2

OpenStudy (amistre64):

these lines meet at: 3x+2 = (-1/3)x +2 3x + 1/3x = 0 10/3x = 0 x = 0

OpenStudy (amistre64):

they appear to meet at the origin...

OpenStudy (amistre64):

they meet at (0,2) if anything lol

OpenStudy (amistre64):

id go with 4

OpenStudy (anonymous):

Per distance formula is \[d= \frac{ \left| ax +by +c \right| }{ \sqrt{a^2 +b^2}}\]

OpenStudy (amistre64):

im prolly wrong since i guessed that lol

OpenStudy (amistre64):

y = 3x+2 ..... 0 = -6/3 +b 0 = -2 +b b = 2 y = (-1/3)x +2.... that seems to work out

OpenStudy (anonymous):

y= 3x+2 get into general form 3x - y + 2 =0 so this means that a=3 , b=-1 and c=2 in our formulas above, and (x,y) are the coordinates of the point ( 6,0) so x=6 , y=0

OpenStudy (anonymous):

d= (9 +2 ) / 2 = 4.5 units

OpenStudy (anonymous):

wait thats wrong

OpenStudy (amistre64):

3x+2 = (-1/3)x +2 3x = -1/3x when x=0 right?

OpenStudy (amistre64):

(0,2) (6,0) 36 + 4 = 40 = sqrt(40)

OpenStudy (amistre64):

2sqrt(10)

OpenStudy (amistre64):

got it ;)

OpenStudy (anonymous):

d = \[\frac { \left| 3(6) + 0 + 2 \right|}{\sqrt{10} }\]

OpenStudy (amistre64):

now your cooking with gas;)

OpenStudy (anonymous):

20/ sqrt (10)

OpenStudy (anonymous):

wtf

OpenStudy (amistre64):

20sqrt(10) --------- 10

OpenStudy (anonymous):

yes I know

OpenStudy (amistre64):

2 sqrt(10)

OpenStudy (amistre64):

need a calculator? :)

OpenStudy (anonymous):

Very puzzling question lol

OpenStudy (amistre64):

just had to get the sleep of my ears to see what i was doin

OpenStudy (anonymous):

Well thank you